# A Fundamental Trade-off in Aligned Language Models and its Relation to Sampling Adaptors

### v1

@@ 7 Results
In both toy and empirical settings, we observe at the corpus level a strong linear anti-correlation between the average log-probability $\frac{\log{p}({\mathcal{Y}})}{N}$ and reward $\frac{{{r}}({\mathcal{Y}})}{N}$. In the toy experiment, as shown in Fig. 1 corpora in the typical set have average log-probabilities and average rewards that exhibit a Pearson correlation of ${r}=-0.92$. And, importantly, these typical corpora occur with significantly higher probability than those that do not. For example, the median log-probability difference between a corpus in and out of the typical set is $10^{84}$ fold.888500,523500523-500,523- 500 , 523 vs. $-500,718$; the difference is somewhat masked by the log-scale. The results in the empirical experiment with real language models are generally consistent with the toy experiment. We observe a Pearson correlation between the average log-probability and average reward of ${r}=-0.93$ with relatively few outliers.
Sampling Adaptors Control the Trade-off.
We observe in Fig. 2 that corpora sampled with different sampling adaptors are centered at different points on the trade-off and follow qualitatively expected trends. For example, corpora sampled with different temperatures have average log-probabilities that follow the expected order —high temperature corpora have lower average log-probabilities and higher average reward. And, at ${\tau}=1.0$ all sampling adaptors produce corpora with higher average log-probability and lower average reward than ancestral sampling. These results are expected since lower temperature and the examined sampling adaptors skew the sampling distribution towards high probability strings. The behaviour when comparing sampling adaptors with different degrees of truncation also follows expectations, e.g., corpora sampled with nucleus sampling for ${\pi}=0.95$ have lower average reward than those sampled with ${\pi}=0.90$. These findings are in line with our theoretical exposition in § 5.2 and suggest that we can use sampling adaptors to control the average reward of sampled corpora.
@@ 8 Conclusion
Our work examines the relationship between probability and reward in sampling from RLHF-tuned language models. We have provided a formal argument and empirical evidence that there exists a trade-off between these two quantities when generating text at scale. Notably, this trade-off exists as a consequence of typicality, is independent of the relationship between reward and probability at the string level, and applies to any conditionally aligned language model, not just those aligned with RLHF. Moreover, we have uncovered a new role of sampling adaptors—the choice of sampling adaptor allows us to select how much likelihood we exchange for reward. This finding presents a new direction of research for improving reward alignment or mitigating reward overfitting in RLHF-tuned models, and the development of sampling adaptors for conditional text generation.
@@
There are three main limitations to our work. First, is that we only conduct empirical analysis for English and Transformer-based language models. Second, we don’t experiment over all sampling adaptors, e.g., we did not consider Mirostat-sampling (Basu et al., 2021) or contrastive search decoding (Su et al., 2022) in our experiments. These choices were made because the theory holds independently of these factors, though further work should consider other model architectures, sampling adaptors and models that span a variety of languages and domains. Finally, we have only examined the probability–quality relationship under the paradigm of RLHF (and equivalently, DPO, as we show in App. E), but not other alignment methods like ORPO (Hong et al., 2024) and KTO (Ethayarajh et al., 2024). We leave those to future work.
@@ Acknowledgments
Appendix A Supplementary proofs for § 5
A.1 Proof of Eq. 9
@@ Appendix A Supplementary proofs for § 5
A.2 Proof of the Probability–Quality trade-off
@@
Appendix B Infinite-Entropy Language Models
@@
Appendix C A Tighter (Chernoff) Bound
In this section, we give a tighter concentration inequality than the (standard) one derived with Chebyshev’s inequality. The inequality displayed in Eq. 9 is weak in the sense that the average right hand size is ${\mathcal{O}}(\frac{1}{N})$—ideally, we desire a concentration inequality that is exponential, i.e., ${\mathcal{O}}(\exp(-cN))$ for some constant $c\in{\mathbbm{R}_{+}}$. To prove such a tighter concentration inequality, we define a class of language models that we term sub-exponential language models. We show that both classical $n$-gram language models as well as modern Transformer-based language models are sub-exponential under our definition. We further show that we can apply the Chernoff–Cramér method to argue that the sample entropy collapses around the mean exponentially quickly.
@@ C.1 Prerequisites
A language model ${p}$ is eos-bounded if there exists $c$ such that ${p}({\textsc{eos}}\mid{\boldsymbol{y}})>c>0$ for all ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$. [^9] We note that for autoregressive language models, though tokens are sampled autoregressively from probability distributions over ${\overline{{{\Sigma}}}}$, the language model is a distribution over ${{{{\Sigma}}^{*}}}$. In other words, ${p}({\boldsymbol{y}}\oplus{\textsc{eos}})\leq(1-c)^{|{\boldsymbol{y}}|}$ for all ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$.
@@
$=$ $-\frac{1}{(1-{\gamma})^{2}}{\mathrm{KL}}(z\,\|\,{p})\leq 0$ (21g)
@@
C.2 A Tighter Concentration Bound
@@
Appendix D Sampling Adaptors and String Probability
In this section our goal is to show that using a sampling adaptor ${\boldsymbol{\alpha}}$ modifies string probability under the prior ${p}({\boldsymbol{y}})$ without modifying the likelihood of alignment with preferences ${p}({A}={\textsc{+}}\mid{\boldsymbol{y}})$. Let us begin by decomposing a sampling adaptor into its constitutent components. Inspired by Meister et al. (2023a), we note that most sampling adaptors can be formulated as the composition of truncation and scaling functions. The truncation function ${C}\colon{{{{\Sigma}}^{*}}}\rightarrow{\mathcal{P}}({\overline{{{\Sigma}}}})$ is a function used to find the set of tokens that meets specified criteria given the prior context, so that tokens deemed likely to lead to undesirable text can have their probability reassigned to other tokens, e.g., to only keep the top-${k}$ tokens. The scaling function ${f}\colon{\mathbbm{R}}\rightarrow{\mathbbm{R}}$ is a simple scaling of the token probability, e.g., scaling the probability by $\frac{1}{{\tau}}$ for some temperature parameter ${\tau}\in{\mathbbm{R}_{+}}$. With these definitions we can express a sampling adaptor ${\boldsymbol{\alpha}}$ as:
${\boldsymbol{\alpha}}\left({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\right)({y})={f}({p}({y}\mid{{\boldsymbol{y}}_{<t}})){\mathbbm{1}}{\{{y}\in{{C}\left({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\right)}\}}$ (31)
That is, given a token distribution ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$, we apply the scaling function to scale token probabilities as needed and then remove tokens according to the truncation function to arrive at the output unnormalized distribution. For instance, we can express nucleus sampling (Holtzman et al., 2020) with:
${f}_{\mathrm{nucleus}}\big{(}({p}({y}\mid{{\boldsymbol{y}}_{<t}})\big{)}$ $={\mathbb{I}}\big{(}{p}({y}\mid{{\boldsymbol{y}}_{<t}})\big{)}$ (32a)
@@ Appendix D Sampling Adaptors and String Probability
where ${\mathbb{I}}$ denotes the identity function and ${\pi}\in{\mathbbm{R}_{+}}$ is a hyperparameter. See § D.1 for more examples. Now, let us consider the probability of a string under the resultant distribution ${{\tilde{{p}}}_{+}}$ when a sampling adaptor ${\boldsymbol{\alpha}}$ is used with an aligned model ${{p}_{{\textsc{+}}}}$. If also apply Bayes’ rule, with some rearrangement we have:
${{\tilde{{p}}}_{+}}({\boldsymbol{y}})$ $={\boldsymbol{\alpha}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}})\prod_{t=1}^{|{\boldsymbol{y}}|}{\boldsymbol{\alpha}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({{y}_{t}})$
$=\frac{\left({\mathbbm{1}}{\{{\textsc{eos}}\in{{C}_{\boldsymbol{y}}}\}}\prod_{t=1}^{|{\boldsymbol{y}}|}{\mathbbm{1}}{\{{{y}_{t}}\in{{C}_{t}}\}}\right)\left({f}\big{(}{p}({\textsc{eos}}\mid{\boldsymbol{y}})\big{)}\prod_{t=1}^{|{\boldsymbol{y}}|}{f}\big{(}{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}},{A}={\textsc{+}})\big{)}\right)}{{{Z}_{\boldsymbol{\alpha}}({\boldsymbol{y}})}}$ (33a)
$\propto\left({\mathbbm{1}}{\{{\textsc{eos}}\in{{C}_{\boldsymbol{y}}}\}}\prod_{t=1}^{|{\boldsymbol{y}}|}{\mathbbm{1}}{\{{{y}_{t}}\in{{C}_{t}}\}}\right)\left({p}({\textsc{eos}}\mid{\boldsymbol{y}})\prod_{t=1}^{|{\boldsymbol{y}}|}{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}}\right){p}({A}={\textsc{+}}\mid{\boldsymbol{y}})$ (33b)
$=\underbrace{{p}({\textsc{eos}}\mid{\boldsymbol{y}}){\mathbbm{1}}{\{{\textsc{eos}}\in{{C}_{\boldsymbol{y}}}\}}\prod_{t=1}^{|{\boldsymbol{y}}|}{p}({\textsc{eos}}\mid{{\boldsymbol{y}}_{<t}}){\mathbbm{1}}{\{{{y}_{t}}\in{{C}_{t}}\}}}_{\text{Modified Prior}}{p}({A}={\textsc{+}}\mid{\boldsymbol{y}})$ (33c)
where we use the shorthands ${{C}_{t}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{{C}\left({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\right)}$ and ${{C}_{\boldsymbol{y}}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{{C}\left({p}(\cdot\mid{\boldsymbol{y}})\right)}$. What Eq. 33 says is that by decomposing a sampling adaptor into its constituent components, we find that we can can factorize out its effects and push them entirely to the prior. As a result, sampling adaptors that can be expressed this way—such as those in § D.1—can be used to control a string’s probability under the prior without affecting its likelihood of alignment with human preferences.
D.1 Examples of Sampling Adaptors

### v2

@@ 7 Results
In both toy and empirical settings, we observe at the corpus level a strong linear anti-correlation between the average log-probability $\frac{\log{p}({\mathcal{Y}})}{N}$ and reward $\frac{{{r}}({\mathcal{Y}})}{N}$. In the toy experiment, as shown in Fig. 1 corpora in the typical set have average log-probabilities and average rewards that exhibit a Pearson correlation of ${r}=-0.92$. And, importantly, these typical corpora occur with significantly higher probability than those that do not. For example, the median log-probability difference between a corpus in and out of the typical set is $10^{84}$ fold.777500,523500523-500,523- 500 , 523 vs. $-500,718$; the difference is somewhat masked by the log-scale. The results in the empirical experiment with real language models are generally consistent with the toy experiment. We observe a Pearson correlation between the average log-probability and average reward of ${r}=-0.93$ with relatively few outliers.
globally normalized Sampling Adaptors Control the Trade-off.
We observe in Fig. 2 that corpora sampled with various globally normalized sampling adaptors are centered at different points on the trade-off and follow qualitatively expected trends. For example, corpora sampled with different temperatures have average log-probabilities that follow the expected order —high temperature corpora have lower average log-probabilities and higher average reward. And, at ${\tau}=1.0$ all sampling adaptors produce corpora with higher average log-probability and lower average reward than ancestral sampling. These results are expected since lower temperature and the examined sampling adaptors skew the sampling distribution towards high probability strings. The behaviour when comparing sampling adaptors with different degrees of truncation also follows expectations, e.g., corpora sampled with nucleus sampling for ${\pi}=0.95$ have lower average reward than those sampled with ${\pi}=0.90$. These findings are in line with our theoretical exposition in § 5.2 and suggest that we can use sampling adaptors to control the average reward of sampled corpora.
Local vs. Global Normalization.
In practice, we find that the tradeoff can be controlled by directly using locally normalized sampling adaptors, i.e., without applying the IMHA post-hoc to derive a globally normalized sampling adaptor. First, we see similar results when we repeat the experiment in § 6.2 without applying the IMHA’s acceptance-rejection protocol. Second, for every locally normalized sampling adaptor over $95\%$ of the samples are accepted by the IMHA. An acceptance rate of $100\%$ implies that the proposal distribution (in this case, the distribution induced by the locally normalized sampling adaptor) is equal to the target distribution (Metropolis et al., 1953; Hastings, 1970; Deonovic and Smith, 2017; Wang, 2022).
@@ 8 Conclusion
Our work examines the relationship between probability and reward in sampling from RLHF-tuned language models. We provide a formal argument and empirical evidence of a trade-off between these two quantities when generating text at scale. Notably, this trade-off exists as a consequence of typicality, is independent of the relationship between reward and probability at the string level, and applies to any conditionally aligned language model, not just those aligned with RLHF.
Moreover, we have proposed globally normalized sampling adaptors, and demonstrate their utility for selecting how much likelihood we exchange for reward. We also find that locally normalized sampling adaptors are good approximations of their globally normalized counterparts in practice, and can be directly used to control the trade-off. Altogether, these findings present a new direction of research for improving reward alignment or mitigating reward overfitting in RLHF-tuned models, and the development of sampling methods for conditional text generation.
@@
There are three main limitations to our work. First, is that we only conduct empirical analysis for English and Transformer-based language models. Second, we don’t experiment over all sampling adaptors, e.g., we did not consider Mirostat-sampling (Basu et al., 2021) or contrastive search decoding (Su et al., 2022) in our experiments. These choices were made because the theory holds independently of these factors, though further work should consider other model architectures, sampling adaptors and models that span a variety of languages and domains. Finally, we have only examined the probability–quality relationship under the paradigm of RLHF (and equivalently, DPO, as we show in App. F), but not other alignment methods like ORPO (Hong et al., 2024) and KTO (Ethayarajh et al., 2024). We leave those to future work.
@@ Acknowledgments
Appendix A Locally-normalized Sampling Adaptors
Despite their empirical success, the normalization performed in locally normalized sampling adaptors can lead to strings that are scored higher by the transform function ${\boldsymbol{\gamma}}$ to have lower probability under the induced language model ${\ddot{{p}}}$. This behaviour stems from the fact that the normalization is dependent on the weights assigned to other tokens at a given time step, and thus leads to inconsistencies at the global level. This is arguably undesirable since it makes it difficult to formally reason about how the transform function—which embodies the core logic of the sampling adaptor—influences the properties of strings sampled from ${\ddot{{p}}}$. To make this clear, we provide an example of this behaviour in top-${k}$ sampling (Fan et al., 2018).
Example 1.
Consider an alphabet ${\overline{{{\Sigma}}}}={\{a,b,c,{\textsc{eos}}\}}$. Let us define a language model [^8] Note that this language model is not necessarily tight. such that
${p}({\boldsymbol{y}})={p}({\textsc{eos}}\mid{\boldsymbol{y}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}})$ (14)
and
${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})=\begin{cases}[0.4,0.1,0.1,0.4]&\text{if ${y}_{t-1}=a$}\\ [0.1,0.4,0.2,0.3]&\text{if ${y}_{t-1}=b$}\\ [0.5,0.5,0.0,0.0]&\text{if $t=1$}\\ \end{cases}$ (15)
where the vector []delimited-[][$\dots$][ … ] denotes the probability assigned to $a,b,c,{\textsc{eos}}$, in that order. Let us now consider the probability of the strings $aa$ and $bb$.
${p}(aa)=0.5\times 0.4\times 0.4\times 0.4=0.032$ (16)
${p}(bb)=0.5\times 0.4\times 0.4\times 0.3=0.024$ (17)
Let us now consider their probability when top-$2$ sampling is applied.
${\ddot{{p}}}(aa)=0.5\times 0.5\times 0.5\times 0.5=0.0625$ (18)
${\ddot{{p}}}(bb)=0.5\times\frac{0.4}{0.7}\times\frac{0.4}{0.7}\times\frac{0.3}{0.7}=0.06997$ (19)
Because the transform function in top-${k}$ does not modify the token probability if the token is kept (see § E.1), Eq. 16 is precisely the score assigned to the strings by the transform function, and Eq. 18 is the score after the normalization step. And we observe a reversal—the string “$aa$” has a higher score than “$bb$”, but is later assigned a lower probability under the induced model ${\ddot{{p}}}$.
Appendix B Supplementary proofs for § 5
B.1 Proof of Eq. 11
@@ Appendix B Supplementary proofs for § 5
B.2 Proof of the Probability–Quality trade-off
@@
Appendix C Infinite-Entropy Language Models
@@
Appendix D A Tighter (Chernoff) Bound
In this section, we give a tighter concentration inequality than the (standard) one derived with Chebyshev’s inequality. The inequality displayed in Eq. 11 is weak in the sense that the average right hand size is ${\mathcal{O}}(\frac{1}{N})$—ideally, we desire a concentration inequality that is exponential, i.e., ${\mathcal{O}}(\exp(-cN))$ for some constant $c\in{\mathbbm{R}_{>0}}$. To prove such a tighter concentration inequality, we define a class of language models that we term sub-exponential language models. We show that both classical $n$-gram language models as well as modern Transformer-based language models are sub-exponential under our definition. We further show that we can apply the Chernoff–Cramér method to argue that the sample entropy collapses around the mean exponentially quickly.
@@ D.1 Prerequisites
A language model ${p}$ is eos-bounded if there exists $c$ such that ${p}({\textsc{eos}}\mid{\boldsymbol{y}})>c>0$ for all ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$. [^9] We note that for autoregressive language models, though tokens are sampled autoregressively from probability distributions over ${\overline{{{\Sigma}}}}$, the language model is a distribution over ${{{{\Sigma}}^{*}}}$.author=ryan,color=violet!40,size=,fancyline,caption=,]Rewrite this footnote. The point is that they are *not* equivalent. The is that ours is exact and theirs is an approximation. I called it “local renormalization”. In other words, ${p}({\boldsymbol{y}}\oplus{\textsc{eos}})\leq(1-c)^{|{\boldsymbol{y}}|}$ for all ${\boldsymbol{y}}\in{{{{\Sigma}}^{*}}}$.
@@
$=$ $-\frac{1}{(1-{\gamma})^{2}}{\mathrm{KL}}(z\,\|\,{p})\leq 0,$ (29g)
@@
D.2 A Tighter Concentration Bound
@@
Appendix E Sampling Adaptors and String Probability
In this section, our goal is to show that an appropriate choice of globally normalized sampling adaptor can be used to control generated strings’ average probability under the prior ${p}({\boldsymbol{y}})$. Let us begin by decomposing a transform function ${\boldsymbol{\gamma}}$ into its constiutent components. Inspired by Meister et al. (2023a), we note that most ${\boldsymbol{\gamma}}$ can be formulated as the composition of truncation and scaling functions. The truncation function ${C}\colon{{{{\Sigma}}^{*}}}\rightarrow{\mathcal{P}}({\overline{{{\Sigma}}}})$ is a function used to find the set of tokens that meets specified criteria given the prior context, so that tokens deemed likely to lead to undesirable text can have their probability reassigned to other tokens, e.g., to only keep the top-${k}$ tokens. The scaling function ${\kappa}\colon{{\mathbbm{R}_{>0}}^{\mid{\overline{{{\Sigma}}}}\mid}}\rightarrow{{\mathbbm{R}_{>0}}^{\mid{\overline{{{\Sigma}}}}\mid}}$ is a simple scaling of the token probability, e.g., to the power of $\frac{1}{{\tau}}$ for some temperature parameter ${\tau}\in{\mathbbm{R}_{>0}}$. With these definitions we can express a transform function ${\boldsymbol{\gamma}}$ as
${\boldsymbol{\gamma}}\left({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\right)({y})={\kappa}({p}({y}\mid{{\boldsymbol{y}}_{<t}})){\mathbbm{1}}{\{{y}\in{{C}\left({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\right)}\}}.$ (39)
That is, given a token distribution ${p}(\cdot\mid{{\boldsymbol{y}}_{<t}})$, we apply the scaling function to scale token probabilities as needed and then remove tokens according to the truncation function to arrive at the output unnormalized distribution. For instance, we can express the transform function in nucleus sampling (Holtzman et al., 2020) with:
${\kappa}_{\mathrm{nucleus}}\big{(}({p}({y}\mid{{\boldsymbol{y}}_{<t}})\big{)}$ $={\mathbb{I}}\big{(}{p}({y}\mid{{\boldsymbol{y}}_{<t}})\big{)}$ (40a)
@@ Appendix E Sampling Adaptors and String Probability
where ${\mathbb{I}}$ denotes the identity function and ${\pi}\in{\mathbbm{R}_{>0}}$ is a hyperparameter. See § E.1 for more examples.
Now, let us consider the probability of a string under the induced distribution ${{\widetilde{{p}}}_{+}}$ when a globally normalized sampling adaptor is used with an aligned model ${{p}_{{\textsc{+}}}}$. If we also apply Bayes’ rule, with some rearrangement we have:
${{\widetilde{{p}}}_{+}}({\boldsymbol{y}})$ $=\frac{{\boldsymbol{\gamma}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({{y}_{t}})}{{{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}}$
$=\frac{\left({\mathbbm{1}}{\{{\textsc{eos}}\in{{C}_{|{\boldsymbol{y}}|}}\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\mathbbm{1}}{\{{{y}_{t}}\in{{C}_{t}}\}}\right)\left({\kappa}\big{(}{p}({\textsc{eos}}\mid{\boldsymbol{y}})\big{)}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\kappa}\big{(}{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}},{A}={\textsc{+}})\big{)}\right)}{{{Z}_{\boldsymbol{\gamma}}({{\Sigma}})}}$ (41a)
$\propto\left({\kappa}_{|{\boldsymbol{y}}|}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\kappa_{t}}\right)\left({\mathbbm{1}}{\{{\textsc{eos}}\in{{C}_{|{\boldsymbol{y}}|}}\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\mathbbm{1}}{\{{{y}_{t}}\in{{C}_{t}}\}}\right)\left({p}({\textsc{eos}}\mid{\boldsymbol{y}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}}\right){p}({A}={\textsc{+}}\mid{\boldsymbol{y}})$ (41b)
$=\underbrace{\left({\kappa}_{|{\boldsymbol{y}}|}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\kappa_{t}}\right)\left({\mathbbm{1}}{\{{\textsc{eos}}\in{{C}_{|{\boldsymbol{y}}|}}\}}{\prod^{|{\boldsymbol{y}}|}_{t=1}}{\mathbbm{1}}{\{{{y}_{t}}\in{{C}_{t}}\}}\right)}_{\text{Functions of the Prior}}{{p}_{{\textsc{+}}}}({\boldsymbol{y}})$ (41c)
where we use the shorthands ${\kappa_{t}}=\frac{{{\kappa}({p}(\cdot\mid{{\boldsymbol{y}}_{<t}}))({{y}_{t}})}}{{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}})}$, ${{C}_{t}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{{C}\left({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\right)}$ and ${{C}_{|{\boldsymbol{y}}|}}\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{{C}\left({p}(\cdot\mid{\boldsymbol{y}})\right)}$. What Eq. 41 says is that by decomposing a transform function ${\boldsymbol{\gamma}}$ into its constitutent components, we see that the resulting probability of a string ${{\widetilde{{p}}}_{+}}({\boldsymbol{y}})$ is entirely determined by the choice of scaling and truncation functions, which take as input the probability of the string under the prior, i.e., ${p}({\boldsymbol{y}})$. As a result, an appropriate choice of globally normalized sampling adaptor can be used to effectively select strings based on their probability under the prior. [^11] So long as the transform functions can be expressed this way—such as those in § E.1. For example, using the transform function from top-${k}$ sampling will lead to the generation of corpora with higher average probability under the prior. [^12] For ${k}<|{\overline{{{\Sigma}}}}|$.
E.1 Examples of Sampling Adaptors